# The Madelung Problem of Finite Crystals

### v1

@@ Abstract
Table 5: Electrostatic energies of ions and a formula unit in a bulk CaTiO3 crystal, where the fraction coordinates are: Ca2+ at (0.5, 0.5, 0.5), Ti4+ at (0, 0, 0), and O2- at (0.5, 0, 0), (0, 0.5, 0), and (0, 0, 0.5).  [^4] Energies are normalized by $e^{2}/d^{2}$, where $d$ is the distance between the nearest Ti4+–O2- pair.
@@
To compute Madelung constants for NaCl, ZnS, CaF2 and CaTiO3, we employ four displacement vectors. One is ${\mathbf{v}}_{1}=(0.5,0.5,0.5)$ as in CsCl; the others (${\mathbf{v}}_{2}$, ${\mathbf{v}}_{3}$, and ${\mathbf{v}}_{4}$) and their corresponding $\nu_{\mathbf{p}bc}$ value are given in Tab. 2. Using Eq. (12), the Madelung constants reported in Tables 4 and 4 are obtained by summing the appropriately weighted $\nu_{\rm pbc}({\mathbf{v}}_{j})$ terms, each analytically corrected via $\nu_{\rm b}$ and $\nu_{\rm corr}$. With $p=20$, nine-digit accuracy is achieved, consistent with the precision of $\nu_{\rm pbc}$ itself. This ${\cal O}(p^{-4})$ EC approach therefore surpasses the CS approach for all cubic lattices investigated.
@@
In summary, we have conducted a systematic analysis of the Madelung problem for finite crystals within a general orthogonal lattice. Our work resolves a conceptual ambiguity by cleanly separating the boundary correction from finite-size corrections in the direct summation. The resulting closed-form boundary correction provides an accurate method for computing Madelung constants with ${\cal O}(p^{-2})$ precision. For cubic lattices, where explicit closed-form expressions for both the boundary and the leading finite-size corrections are available, the approach achieves ${\cal O}(p^{-4})$ accuracy. Even with minimal supercells—corresponding to $p=1$ (i.e., $3^{3}$ unit cells)—practical calculations already attain accuracy on the order of $10^{-4}$, demonstrating applicability to ionic crystals such as NaCl, ZnS, CaF2 and CaTiO3.
This work was supported by NSFC (Grant Nos. 22273047 and 21873037).
@@
.2 Derivation of $I_{1}$ via the divergence theorem
@@
.3 Derivation of $I_{1}$ via the Fourier transform
@@
This alternative derivation has certain advantanges. For example, from Eqs. (S30), (S31) and the Dirichlet integral
@@
The infinite boundary term $\nu_{\rm ib}(\mathbf{r})=I_{1}({\mathbf{r}})/V$ is directly related to $J({\mathbf{M}})$ via the following pairwise sumHu (2014); Yi et al. (2017)
$\sum_{i<j}^{N}q_{i}q_{j}\nu_{\rm ib}({\mathbf{r}}_{i}-{\mathbf{r}}_{j})=J({\mathbf{M}})-\frac{q_{\rm tot}}{V}\sum_{j=1}^{N}q_{j}|{\mathbf{r}}_{j}|^{2},$ (S41)
@@
.4 Derivation of $I_{2}$ via the divergence theorem

### v2

@@ V Applications to various lattices
Table 5: Electrostatic energies of ions and a formula unit in a bulk CaTiO3 crystal, where the fraction coordinates are: Ca2+ at (0.5, 0.5, 0.5), Ti4+ at (0, 0, 0), and O2- at (0.5, 0, 0), (0, 0.5, 0), and (0, 0, 0.5), Energies are normalized by $e^{2}/d^{2}$, where $d$ is the distance between the nearest Ti4+–O2- pair.
@@
To compute Madelung constants for NaCl, ZnS, CaF2 and CaTiO3, we employ four displacement vectors. One corresponds to the CsCl structure, given by ${\mathbf{v}}_{1}=(0.5,0.5,0.5)$; the remaining vectors (${\mathbf{v}}_{2}$, ${\mathbf{v}}_{3}$, and ${\mathbf{v}}_{4}$) and their associated $\nu_{\rm pbc}$ are listed in Tab. 2. Using Eq. (27), we obtain the Madelung constants presented in Tabs. 4 and 5. These constants are calculated by summing the appropriately weighted $\nu_{\rm pbc}({\mathbf{v}}_{j})$ terms, with each term analytically corrected via $\nu_{\rm b}$ and $\nu_{\rm corr}$. Our computed reference values for NaCl and ZnS agree with established literature datanac ; zns . For CaF2, the Ca2+ ion occupies an environment identical to that of either Zn2+ or S2- in the ZnS structure. After multiplying the sum of the Madelung constants for Ca2+ and F- by $2$, the result also aligns with published valuescaf . Notably, with $p=20$, we achieve nine-digit accuracy, which matches the intrinsic precision of $\nu_{\rm pbc}$ itself. This ${\cal O}(p^{-4})$ EC approach therefore surpasses the CS approach for all cubic lattices examined.
@@
VI Conclusions
In summary, we have conducted a systematic analysis of the Madelung problem for finite crystals within a general orthogonal lattice. Our work resolves a conceptual ambiguity by cleanly separating the boundary correction from finite-size corrections in the direct summation. The resulting closed-form boundary correction provides an accurate method for computing Madelung constants with ${\cal O}(p^{-2})$ precision. For cubic lattices, where explicit closed-form expressions for both the boundary and the leading finite-size corrections are available, the approach achieves ${\cal O}(p^{-4})$ accuracy. Even with minimal supercells—corresponding to $p=1$ (i.e., $3^{3}$ unit cells)—practical calculations already attain accuracy on the order of $10^{-4}$, demonstrating applicability to ionic crystals such as CsCl, NaCl, ZnS, CaF2 and CaTiO3.
This work was supported by NSFC (Grant Nos. 22273047 and 21873037). Z.H is grateful to Andrei Zabolotskii for pointing out relevant references to Madelung constants.
Supporting Information: complete derivations of the boundary and finite-size corrections [Eqs. (3) and (32)].
VII Appendix: Derivation of the Indefinite Integral via Euler Substitution
In this appendix, we evaluate the following integral:
$I=\int\frac{du}{\left(u^{2}+1\right)\sqrt{u^{2}+2}}.$ (A1)
Using the Euler substitution
$\sqrt{u^{2}+2}+u=\tau,$ (A2)
which maps the domain $u\in(-\infty,\infty)$ to $\tau\in(0,\infty)$, we obtain the relations:
$\sqrt{u^{2}+2}=\frac{\tau}{2}+\frac{1}{\tau}=\tau\,\frac{du}{d\tau},$ (A3)
and
$u^{2}+1=\frac{\tau^{2}}{4}+\frac{1}{\tau^{2}}.$ (A4)
Substituting these expressions into the integral yields
$\begin{aligned} I&=\int d\tau\frac{4\tau}{\tau^{4}+4}={\rm atan}\frac{t^{2}}{2}+C\\ &={\rm atan}\frac{\left(\sqrt{u^{2}+2}+u\right)^{2}}{2}+C\end{aligned},$ (A5)
where $C$ is an integration constant and ${\rm atan}(\cdot)$ denotes the inverse tangent function . Furthermore, applying the subtraction formula
${\rm atan}A-{\rm atan}B={\rm atan}\frac{A-B}{1+AB},$ (A6)
valid when AB>1AB>-1, we find
${\rm atan}\frac{\left(\sqrt{u^{2}+2}+u\right)^{2}}{2}-{\rm atan}\frac{u}{\sqrt{u^{2}+2}}=\frac{\pi}{4}.$ (A7)
Consequently, the integral can be expressed in the alternative form
$I={\rm atan}\frac{u}{\sqrt{u^{2}+2}}+C^{\prime},$ (A8)
where the constant $C^{\prime}$ differs form $C$ by $\pi/4$. The Euler substitution and the arctangent identity presented here are also utilized to derive analytical results in the Supporting Information. In the limit $p\to\infty$, the difference between the cuboid and cubic crystals results in an additional electrostatic potential at the origin, given by
$\int_{1}^{3}\frac{du\,8t^{2}}{\left(u^{2}+1\right)\sqrt{u^{2}+2}}=8t^{2}\left({\rm atan}\frac{3}{\sqrt{11}}-\frac{\pi}{6}\right)\\ =4t^{2}\left(\frac{\pi}{6}-{\rm atan}\frac{1}{3\sqrt{11}}\right),$ (A9)
where the final expression follows directly from the difference between the two boundary terms (see Eq. (32)).
@@ VII.1 Preliminary Integrals
VII.2 Derivation of $I_{1}$ via the divergence theorem
@@
VII.3 Derivation of $I_{1}$ via the Fourier transform
@@
This alternative derivation has certain advantages. For example, from Eqs. (S30), (S31) and the Dirichlet integral
@@
The infinite boundary term $\nu_{\rm b}(\mathbf{r})=I_{1}({\mathbf{r}})/V$ is directly related to $J({\mathbf{M}})$ via the following pairwise sumHu (2014a); Yi et al. (2017)
$\sum_{i<j}^{N}q_{i}q_{j}\nu_{\rm b}({\mathbf{r}}_{i}-{\mathbf{r}}_{j})=J({\mathbf{M}})-\frac{q_{\rm tot}}{V}\sum_{j=1}^{N}q_{j}|{\mathbf{r}}_{j}|^{2},$ (S41)
@@
VII.4 Derivation of $I_{2}$ via the divergence theorem